40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 28 02:33:04 2011, Jenne, Update, Locking, Yarm flashing, but not yet locked 
    Reply  Thu Jul 28 16:10:34 2011, kiwamu, Update, LSC, Both arm locked 
Message ID: 5054     Entry time: Thu Jul 28 16:10:34 2011     In reply to: 5051
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Both arm locked 

[Nicole / Jamie / Rana / Kiwamu]

  The X arm and Y arm have been locked.

The settings for the locking were stored on the usual IFO_CONFIGURE scripts, so anybody can lock the arms.

In addition to that Nicole, Jamie and Rana re-centered the beam spot on the ETMY_TRANS camera and the TRY PD.

The next step is to activate the C1ASS servo and align the both arms and beam axis.

 

 

Xarm locking notes:

* Changed TRX gain from -1 to -0.02. Without this 50x reduction the arm power was not normalized.

* Had to fix trigger matrix to use TRX for XARM and TRY for YARM. Before it was crazy and senseless.

* Lots of PZT alignment. It was off by lots.

* Yarm trans beam was clipping on the steering mirrors. Re-aligned. Needs to be iterated again. Be careful when bumping around the ETMY table.

* YARM gain was set to -2 instead of -0.2. Because the gain was too high the alignment didn't work right.

ALWAYS HAVE an OPEN DATAVIEWER with the standard ARM channels going when doing ANY INTERFEROMETER WORK.

THIS IS THE LAW.

ELOG V3.1.3-